# Anyview: Generalizable Indoor 3D Object Detection with Variable Frames

### v1

@@ V-B Results and Analysis
On wider range of input views: We further extend ScanNet-MV to a series of fine-grained benchmarks for more comprehensive analysis. We train models on 10/30/50 views uniformly sampled from the whole 50 views in ScanNet-MV and evaluate them on different numbers of views. We devise two settings for partial scenes: Uniform and Continuous, indicating the evaluation of model is conducted on uniformly/adjacently sampled views. As shown in Table IV, 3DETR-m trained on small number of views gets low performance on large number of views, and vice versa. On the contrary, with the growing of numbers of training views, AnyView shows consistent improvement across various scales of input. Even if combining the best results of three 3DETR-m models, AnyView still surpasses them by a large margin with a single set of parameters.
@@ V-C Ablation Study
We conduct ablation studies to show how different architecture designs and training/inference strategies influence the performance of the proposed AnyView framework. The models are trained on ScanNet-MV with 30 views.
[…]
TABLE V: The effects of global query sampling, coordinate transformation and positional embeddings for transformer encoder on the final performance of AnyView on ScanNet-MV.
[…]
@@
Figure 5: Online detection visualization results. We demonstrate the detection results for 15, 30, 40, and 50 frames during the navigation(scene0086_00, scene00430_00 and scene0474_00 on the ScanNet dataset).Red bounding boxes and green bounding boxes represent false-positive and true-positive samples, respectively.
@@
V-D Online 3D Object Detection
We conduct online object detection experiments to demonstrate the performance of AnyView with any-scale point cloud input. We select 3DETR as the baseline network, and these models are all trained on ScanNet-MV with 50 views. During the validation process, the views of each scene are fed into AnyView in a frame-by-frame format, and the aggregated point cloud reconstruction results are fed into the 3DETR.
Quantitative results: We demonstrate the average accuracy of AnyView and 3DETR in the online detection setting, as illustrated in Table VI. Our proposed AnyView achieves an average accuracy of (55.32/30.94) in the online detection setting. Compared to the MV(D) setting, which only decreases by 5.47 and 4.86 on the mAP@25 and mAP@50 measures, respectively. However, since 3DETR is unable to adapt to any size point cloud, the accuracy decreases significantly in the online setting, by 13.61 vs. 11.48, respectively. Meanwhile, we have visualized the curve of model performance with frame rate, as shown in Fig. 7. As the number of input RGB-D frames is too few, both AnyView and 3DETR are less accurate due to the sparse point cloud, though AnyView still outperforms 3DETR by 7.26 vs. 6.68. As the number of input RGB-D frames increases, AnyView improves accuracy at a faster rate and stabilizes (View Number = 10). Meanwhile, the accuracy of AnyView has reached 52.78 vs. 29.18, which is only 6.71 vs. 6.64 lower than the overall scene. However, the accuracy of 3DETR has been improving with the number of reconstructed scene frames due to the failure to adapt to sparsely distributed scene point clouds. The performance of 3DETR stabilizes only after 40 frames of RGB-D image reconstruction and is 4.62 lower on the mAP@25 measure compared to AnyView. All the above results indicate that our proposed AnyView can be able to handle scene point clouds of any size and has the ability of online detection.
Qualitative results: We visualize the detection results of AnyView with 3DETR at different frame numbers as illustrated in Fig. 5. In the case of a large degree of site cloud fragmentation (15 Views), 3DETR has more false-positive samples, while AnyView still maintains a high detection accuracy. As the number of viewpoints of the reconstructed scene point cloud increases, the false-positive samples detected by 3DETR are suppressed to some extent, and better results are achieved at 50 Views. While the detection results of AnyView remain stable, the increase in the number of viewpoints does not lead to a significant increase in performance. The best performance is also achieved at 50 Views. The visualization results of online detection further demonstrate that the proposed AnyView can efficiently handle input scenes with different numbers of views and has the capability of online detection.
Figure 7: The effect of the number of image frames on the overall scene detection accuracy in the Online setting.
TABLE VI: 3D object detection results at the online setting (MAP@0.25 and MAP@0.5), where S-L denotes Scene-level
Method Training Evaluation mAP
Benchmark Benchmark @0.25 @0.5
S-L 3DETR MV(D) Online 41.09 23.82
AnyView MV(D) Online 55.23 30.94
@@ VI Conclusion
In this paper, we challenge the existing benchmarks for indoor 3D object detection in a perspective of input data. We propose a new practical setting for this task, which unifies the input modality as multi-view RGB-D images with variable input frame numbers and evaluates 3D detectors on various scales of input data. We design a new transformer-based framework named AnyView for practical applications, which is able to process scenes consisting of any number of frames and extract input scale-independent scene representations. Benefiting from the nature of the self-attention mechanism and our scale-independent training strategy, AnyView is able to change the number of scene representations extracted for each input frame during inference and flexibly handle various numbers of input frames. Extensive experiments on the ScanNet dataset show that AnyView achieves both great generalizability and high detection accuracy while containing a similar amount of parameters with the baselines.
The main limitation of this work is the weak generalization to new categories of objects in practical scenes. Meanwhile, the dynamic token mechanism is only adaptive in terms of quantity, which may lead to loss of information. Future work hopes to achieve dynamic adaptation with token pruning techniques and open vocabulary detection.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62376032, Grant U22B2050, and Grant 62125603.

### v2

@@ V-B Results and Analysis
Figure 4: Online detection visualization results (IoU=0.25). We demonstrate the detection results for 15, 30, 40, and 50 frames during the navigation(scene0474_00 on the ScanNet dataset).Red bounding boxes and green bounding boxes represent false-positive and true-positive samples, respectively.
On wider range of input views: We further extend ScanNet-MV to a series of fine-grained benchmarks for more comprehensive analysis. We train models on 10/30/50 views uniformly sampled from the whole 50 views in ScanNet-MV and evaluate them on different numbers of views. We devise two settings for partial scenes: Uniform and Continuous, indicating the evaluation of model is conducted on uniformly/adjacently sampled views. As shown in Table IV, 3DETR-m trained on small number of views gets low performance on large number of views, and vice versa. On the contrary, with the growing of number of training views, AnyView shows consistent improvement across various scales of input. Even if combining the best results of three 3DETR-m models, AnyView still surpasses them by a large margin with a single set of parameters.
@@
Online 3D object detection: We conduct online object detection experiments to demonstrate the performance of AnyView with any-scale point cloud input as shown in Fig. 4. The views of each scene are fed into AnyView in a frame-by-frame format, and the aggregated point cloud reconstruction results are fed into the 3DETR.
@@ V-C Ablation Study
We conduct ablation studies to show how different architecture designs and training/inference strategies influence the performance of the proposed AnyView framework. The models are trained on ScanNet-MV with 30 views. With a high degree of site cloud fragmentation (15 views), 3DETR exhibits more false positives, while AnyView maintains high detection accuracy. As the number of viewpoints increases, 3DETR’s false positives are somewhat reduced, achieving better results at 50 views.
Performance w.r.t. architecture design: We ablate three parts of architecture designs of AnyView, as shown in Table V. GlobalQuery indicates the object queries are sampled from the concatenated scene (✓) instead of the coordinates of scene proxies. CoordsTrans means whether to apply geometry learner in the camera coordinate (✓) of each view or in the world coordinate. PEenc indicates whether to use positional embeddings for the transformer encoder. The first two rows show that high object query coverage of the scene is beneficial for the transformer decoder. The third row shows unifying the coordinate makes the geometry learner extract more robust scene proxies. As the features of point clouds already contain positional information, 3DETR finds positional embeddings are not necessary. However, in AnyView the scene proxies are independently extracted, so the spatial relationship between scene proxies from different views is weakened. We find Fourier positional embeddings with a MLP performs best.
TABLE V: The effects of global query sampling, coordinate transformation and encoder positional embeddings on the AnyView with ScanNet-MV.
GlobalQuery CoordsTrans PEenc mAP
@@ VI Conclusion
In this paper, we challenge the existing benchmarks for indoor 3D object detection in a perspective of input data. We propose a new practical setting for this task, which unifies the input modality as multi-view RGB-D images with variable input frame numbers and evaluates 3D detectors on various scales of input data. We design a new transformer-based framework named AnyView for practical applications, which is able to process scenes consisting of any number of frames and extract input scale-independent scene representations. Benefiting from the nature of the self-attention mechanism and our scale-independent training strategy, AnyView is able to change the number of scene representations extracted for each input frame during inference and flexibly handle various numbers of input frames. Extensive experiments show that AnyView achieves both great generalizability and high detection accuracy.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62376032, Grant 62125603 and A*STAR National Robotics Programme (RIE2025) M25N4N2009.